OpenStudy (anonymous):

Help on checking my answer? Indicate the equation of the given line in standard form. The line containing the longer diagonal of a quadrilateral whose vertices are A (2, 2), B(-2, -2), C(1, -1), and D(6, 4). i got -x + y + 6 = 0

OpenStudy (anonymous):

i know the longer diagonals are B to D

OpenStudy (anonymous):

and slope is 3/4

OpenStudy (anonymous):

or x-y+6=0..

OpenStudy (anonymous):

@apoorvk?

OpenStudy (anonymous):

@dpaInc?

OpenStudy (apoorvk):

what equation did you use? your slope is right, but seems like a calculation mistake while finding the line.

OpenStudy (lgbasallote):

@lgbasallote

OpenStudy (anonymous):

y-y1=m(x-x1) y+2=3/4(x+2)

OpenStudy (anonymous):

4y + 8 = 4x + 2 4y + 6 = 4x y + 6 = x

OpenStudy (anonymous):

is that right so far? anyone??

OpenStudy (anonymous):

your fist line, right side is incorrect...

OpenStudy (anonymous):

in the line 4y + 8 = 4x + 2, the right side is incorrect.

OpenStudy (anonymous):

so the left side isnt.

OpenStudy (anonymous):

4y + 8 is correct... i'm assuming you cleared the fraction by multiplyting the whole equation by 4.... that'll leave a 3(x+2) on the right...

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

woops i read that wrong. okay so 4y+8=3x+2 4y+6=3x

OpenStudy (anonymous):

the right side is still wrong .... distribute the 3 into the binomial...

OpenStudy (anonymous):

4y+8=3x+6 4y+2=3x

OpenStudy (anonymous):

whewww....:) ok... proceed to make the equation in general form... make 1 side equal to zero...

OpenStudy (anonymous):

haha sorry, silly mistakes. i'm brain fried. 4y + 2 = 3x -3x + 4y + 2 = 0

OpenStudy (anonymous):

that's good... of course you know you could also write an equivalent equation in: 3x - 4y - 2 = 0 that too, would be correct...

OpenStudy (anonymous):

:) thank you!

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

this 3x - 4y = 2 is still correct, yes?

OpenStudy (anonymous):

yep...

OpenStudy (anonymous):

k